OpenStudy (anonymous):

help solving linear equations and inequalities? I'll medal and fan! Thanks! ♥

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

?? what's ur queastion

OpenStudy (anonymous):

5x+4x+8+7+11x=55 can you help me with that and explain please? thanks :)

OpenStudy (anonymous):

okay that's easy one

OpenStudy (anonymous):

yeah but it gets harder lol

OpenStudy (anonymous):

do I first collect the like terms?

OpenStudy (anonymous):

20x+15=55 20x=40 then x=40/20=2

OpenStudy (anonymous):

yeah exactly

OpenStudy (anonymous):

did u get it ??

OpenStudy (anonymous):

not really haha can you explain?

OpenStudy (anonymous):

okay it simple as it's just collect the x's terms and nubers term

OpenStudy (anonymous):

and then take the divison

OpenStudy (anonymous):

of operand of x

OpenStudy (anonymous):

so the first step is (5x+4x+11x)+(8+7)=55, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then it's obviously 20x+15=55

OpenStudy (anonymous):

then subtract 15 from both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no eaxtly

OpenStudy (anonymous):

so it cancels out the 15 on one side and the other side is =40?

OpenStudy (anonymous):

u just move it out 15 to other side

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

finally divide the operand of x

OpenStudy (anonymous):

ok then divided by 20 so it's x=2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are excellenet

OpenStudy (anonymous):

cool so I have a few more if that's okay with you?

OpenStudy (anonymous):

haha thanks I try :p

OpenStudy (anonymous):

don't forget a medal please

OpenStudy (anonymous):

I won't haha but can you help with a few more?

OpenStudy (anonymous):

why not i'm just helped u

OpenStudy (anonymous):

no I'll medal don't worry lol

OpenStudy (anonymous):

but I still need help if you're available?

OpenStudy (anonymous):

of course if there r more medal

OpenStudy (anonymous):

I'll fan?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright so I think I have the right answer to the next one but I just want to double check it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so it's \[7+4+(3x+x+2x)=9\]

OpenStudy (anonymous):

and I got \[16x=9\]

OpenStudy (anonymous):

6x+11=9 6x=-2 x=-1/3

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

am i get a medal plz

OpenStudy (anonymous):

I already gave you one haha

OpenStudy (anonymous):

plz another one

OpenStudy (anonymous):

I think you can only give one?